How much greater is the boiling point(-29) then melting point(-97)
I did this on a test and i continuisly got it wrong,at first i worked it out like this...
-97-(-29) I changed the - between 97 and 29 to a + (that is what my teacher said to do) and i added 29+97 and ended up with 126,of corse I knew that it was still a negitive number because -29 is larger then -97 so I ended up with
-126....
PLEASE HELP ME IM AM SO CONFUSED!!!!!!!! ANYONE PLEASE!!!!!!!

Answer by Alan3354(69443)

To get to -29 from -97, you move 68 units.
-29 is 68 more than -97.
-----------
-29 - (-97) = -29 + 97 = 68